# Fast Gradient Computation for RoPE Attention in Almost Linear Time

### v1

@@ 1 Introduction
The GPT-o3 [56], Llama 3.3 [53, 3], Claude 3.5 [4] are transformed-based Large Language Models (LLMs), have become important tools in natural language processing, which enables applications from machine translation to sentiment analysis. In the Transformer architecture, attention mechanisms, computationally intensive operations, compute token correlations within the sequence [69]. The efficiency of attention computations, both in forward computations and backward gradient computations, directly influenced the scalability and feasibility of training LLMs, especially when the size and input context length of these LLMs continue to grow [9, 6]. In recent research, Rotating Position Embedding (RoPE) [57] has become a popular modification to the attention mechanism, and it has enabled models to capture positional relationships between tokens with better expressiveness. The RoPE mechanism has been adopted in state-of-the-art models, such as Llama [67, 68, 53], Claude [4], Apple’s LLMs [29, 54], and many others, but the implementation of RoPE complicates attention computation due to the additional structure imposed by position-dependent rotations [57]. In recent work, [7] has demonstrated an efficient algorithm for forward computation of RoPE attention in the bounded entry regime, but the backward computation, the process of calculating gradients for model optimization, is less explored.
@@
To the best of our knowledge, this is the first work to characterize the fine-grained complexity of backward computations in RoPE attentions, extending prior results on forward computations in RoPE attention [7].
@@
In Section 2, we present some relevant papers. In Section 3, we show important components for the RoPE attention. In Section 4, we present the details of getting the gradients of RoPE-based attention mechanisms. Section 5 discusses the exact time complexity for computing RoPE attention gradients. Section 6 shows the fast computation of gradient using the polynomial method. Section 7 details the lower bounds of hardness. Finally, Section 8 provides conclusions and avenues for future work.
@@ 2 Related Work
In this paper, we study a variant of attention called RoPE attention. At a high level, RoPE gives more expressive power to the model in exchange for making the computational problem more complicated. In particular, many prior algorithms, such as the algorithm of [6], no longer apply to RoPE for fundamental reasons we will discuss. RoPE was proposed by [57] and has been used extensively in large-scale industrial models. Examples which are known to use RoPE include the open-source models released by Meta such as Llama [67] (see page 3), Llama 2 [68] (see page 5), Llama 3 [53] (see page 7), and the close-source LLM Claude 3.5 [4] released by Anthropic. Apple also incorporates RoPE into their LLM architecture (see [54], and page 3 of  [29]). The idea behind RoPE is to rotate the query and key vectors in the self-attention mechanism. The rotation is position-dependent and designed such that the inner product between two position-encoded vectors reflects their relative position in the input so that in the RoPE attention mechanism, pairs of tokens with a longer relative distance will have a smaller correlation.
@@
The attention mechanism has often been criticized for its quadratic computational complexity concerning context length, a challenge that becomes more pronounced as the sequence length grows in today’s LLMs [2, 56, 53, 3, 5, 4]. However, this issue can be addressed using polynomial kernel approximation methods [1], which facilitate constructing the approximated attention matrix using low-rank approximations. Such methods enable substantial improvements in computation speed, allowing a single attention layer to perform both training and inference nearly as fast as linear time [6, 8]. Our research further extends this efficiency to support multi-layer transformer architectures for both training and inference. In addition, these techniques can generalize to advanced attention mechanisms, such as tensor attention, while preserving the almost linear time complexity in both training and evaluation phases [9, 50]. Beyond this, alternative theoretical methods also exist. For example, the conv-basis approach introduced in [44] offers another avenue for speeding up attention computation.
@@
The attention has become a cornerstone in AI, particularly in large language models (LLMs), which excel in NLP tasks such as machine translation, text generation, and sentiment analysis due to their ability to capture complex contextual relationships. However, understanding the attention mechanism from a theoretical perspective remains an ongoing challenge. Several works have explored the theoretical foundations and computational complexities of attention [66, 77, 17, 6, 63, 18, 20, 55, 65, 12, 22, 16, 62, 10], focusing on areas such as efficient attention [31, 58, 64, 40, 44, 50, 48, 46, 45, 21, 43, 35, 34, 30, 70, 32, 71, 36, 8, 24, 58, 41, 49], optimization [23, 51, 59, 27, 52, 26, 25], and the analysis of emergent abilities [15, 74, 11, 37, 76, 42, 75, 19, 47, 33, 73, 60, 72, 75, 38]. Notably, [77] introduced an algorithm with provable guarantees for attention approximation, [39] proved a lower bound for attention computation based on the Strong Exponential Time Hypothesis, and [6] provided both an algorithm and hardness results for static attention computation using low-rank approximation [28, 61].
@@ 3.1 Notation
For $n\in\mathbb{Z}^{+}\cup\{0\}$, for set $\{1,2,\cdots,n\}$, we denote the set by using the notation $[n]$. Here, we define the concept of nearly linear time when the time is $O(n\log n)$. We introduce the concept of almost linear time when time is $O(n^{1+o(1)})$. Given $a$ as any vector, we say the diagonal matrix of $c$ is $\operatorname{diag}(c)$ where $c_{i}$ means the $i,i$-th entry in the matrix $\operatorname{diag}(c)$. For any matrix, we denote the support of the matrix using the notation $\operatorname{supp}$, that is, the set of entries where the matrix is nonzero. $B^{\top}$ is defined as $(B^{\top})_{i,j}:=B_{j,i}$. Suppose there are two vectors $c,d$ of the same length. We denote the entry-wise multiplication by using the notation $c\circ d$; that is, the $i$-th entry in that vector is $c_{i}d_{i}$. To denote the Frobenius norm, for any matrix $B$, we denote it as $\|B\|_{F}:=\sqrt{\sum_{i,j}B_{i,j}^{2}}$; to denote the maximum norm of matrix $B$, we use $\|B\|_{\infty}:=\max_{i,j}|B_{i,j}|$. Suppose there are two matrices $C,D$ of the same dimensions. We represent the Hadamard product or the entry-wise multiplication by using the notation $C\circ D$, that is, $(i,j)$-th entry of the matrix is $C_{i,j}\cdot D_{i,j}$. Let $C\in\mathbb{R}^{n_{0}\times m_{0}}$ and $D\in\mathbb{R}^{n_{1}\times m_{1}}$. We define $C\otimes D$ is an $n_{0}n_{1}\times m_{0}m_{1}$ matrix, where $(C\otimes D)_{(j_{0}-1)n_{1}+j_{1},(i_{0}-1)m_{2}+i_{1}}$ is equal to $C_{j_{0},i_{0}}D_{j_{1},i_{1}}$ for any $j_{0}\in[n_{0}],i_{0}\in[m_{0}],j_{1}\in[n_{1}],i_{1}\in[m_{1}]$.
@@ 3.2 Problem Definition
$A_{i,j}(\mathsf{X})=\exp(\underbrace{\mathsf{A}_{i+(j-1)n,*}}_{1\times d^{2}}\underbrace{\mathsf{X}}_{d^{2}\times d^{2}}\underbrace{w_{i-j}/d}_{d^{2}\times 1}).$
@@
$A_{i,j}(\mathsf{X})=$ $~{}\exp(\underbrace{(\mathsf{A}_{i+(j-1)n,*}\otimes w_{i-j}^{\top})}_{1\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1})=~{}\exp(\underbrace{(\mathsf{A}_{i+(j-1)n,*}\otimes\mathsf{W}_{i+(j-1)n,*})}_{1\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1}).$
@@
Then, the optimization problem in the context of RoPE attention computation is described as:
@@
Let $B>0$ and $\epsilon>0$ denote two parameters. Given a set of matrices $W_{-(n-1)},\cdots,W_{-1}$, $W_{0},W_{1},\cdots,W_{n-1}\in\mathbb{R}^{d\times d}$ where $\mathrm{supp}(W_{i})\subset S$ for all $i\in\{-(n-1),\cdots,-1,$ $0,1,\cdots,n-1\}$. Here $S\subseteq[d]\times[d]$ where $|S|=O(d)$. For $i,j\in[n]$, let $\mathsf{W}\in\mathbb{R}^{n^{2}\times d^{2}}$ such that $\mathsf{W}_{i+(j-1)n,*}=\operatorname{vec}({W_{i-j}})$. Here, we suppose four $n\times d$ matrices $A_{1},A_{2},A_{3},E$, and we have three $d\times d$ matrices $X_{1},X_{2},Y$. Let $\mathsf{X}:=X_{1}\otimes X_{2}\in\mathbb{R}^{d^{2}\times d^{2}}$. We define the matrix $A(\mathsf{X})\in\mathbb{R}^{n\times n}$ as the matrix representation of $\exp(\underbrace{(\mathsf{A}\oslash\mathsf{W})}_{n^{2}\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1})$ and the $n\times n$ matrix $D(\mathsf{X}):=\operatorname{diag}(\underbrace{A(\mathsf{X})}_{n\times n}\underbrace{{\bf 1}_{n}}_{n\times 1}).$ The RoPE attention optimization problem is formulated as:
@@ 3.4 Time Complexity of Multiplications
Here, we introduce the time complexity of multiplying matrices as follows:
@@
Based on [13, 14], we have the following fact.
@@ 3.6 Basic Facts
In this section, we introduce some basic facts we will use.
Fact 3.8.
We suppose there are three vectors of $n$ dimension $x,y,z$. Thus, we get
  • $\langle x\circ y,z\rangle=x^{\top}\operatorname{diag}(y)z$.
  • $\langle x,y\rangle=\langle x\circ y,{\bf 1}_{n}\rangle$.
Fact 3.9 (Folklore).
@@
Fact 3.10.
@@
Here, we introduce a technique called the tensor trick.
Fact 3.11 (Tensor trick).
@@
Now, we introduce the following properties.
Fact 3.12.
@@
From Lemma and Def. 3.11, we are able to prove this fact. ∎
@@ 3.7 Useful Definitions
Here, we denote $d^{4}$-dimensional vector $x\in\mathbb{R}^{d^{4}}$ as the vectorization of a $d^{2}\times d^{2}$ matrix $\mathsf{X}$. We introduce some terms to simplify our notations in calculation.
Definition 3.13.
@@
Definition 3.14.
@@
Definition 3.15.
From Def. 3.13, it defines $u(\cdot)_{j_{0}}$ , and we have $\alpha(\cdot)_{j_{0}}$based on Def. 3.14. The function $s(x)_{j_{0}}$ maps a $d^{4}$-dimensional vector to an $n$-dimensional vector given every $j_{0}\in[n]$ such that
@@
Definition 3.16.
@@
From Def. 3.15, with every $j_{0}$ in the set $[n]$, it gives $s(x)_{j_{0}}$ as an $n$-dimensional normalized vector, and we define $v(y)_{i_{0}}$ based on Def. 3.16 given that each $i_{0}\in[d]$. Defining a function $\ell(x)_{j_{0},i_{0}}$ maps a $d^{4}$-dimensional vector to a scalar with each $j_{0}\in[n]$ and each $i_{0}\in[d]$ such that
@@ 3.8 Reformulation of the Loss Function
Based on the notation in Section 3.7, we can reformulate and simplify our loss function.
@@
  • Based on Def. 3.18, For $j_{0}$ in the set $[n]$ and $i_{0}$ in the set[d]$[d]$, we get $\mathsf{Loss}(\mathsf{X})_{j_{0},i_{0}}$
@@
where the 1st equality is based on Def. 3.2, the definition of Frobenius norm derives the 2nd equality, the 3rd equality is due to Def. 3.15 and Def. 3.16, and the 4th step is based on Def. 3.18. ∎

### v2

@@ 1 Introduction
The GPT-o3 [57], Llama 3.3 [54, 3], Claude 3.5 [4] are transformed-based Large Language Models (LLMs), have become important tools in natural language processing, which enables applications from machine translation to sentiment analysis. In the Transformer architecture, attention mechanisms, computationally intensive operations, compute token correlations within the sequence [72]. The efficiency of attention computations, both in forward computations and backward gradient computations, directly influenced the scalability and feasibility of training LLMs, especially when the size and input context length of these LLMs continue to grow [9, 6]. In recent research, Rotating Position Embedding (RoPE) [58] has become a popular modification to the attention mechanism, and it has enabled models to capture positional relationships between tokens with better expressiveness. The RoPE mechanism has been adopted in state-of-the-art models, such as Llama [70, 71, 54], Claude [4], Apple’s LLMs [29, 55], and many others, but the implementation of RoPE complicates attention computation due to the additional structure imposed by position-dependent rotations [58]. In recent work, [7] has demonstrated an efficient algorithm for forward computation of RoPE attention in the bounded entry regime. However, backward computation, the process of calculating gradients for model optimization, has been explored less.
@@
To the best of our knowledge, this is the first work to characterize the fine-grained complexity of backward computations in RoPE attentions, extending prior results on forward computations in RoPE attention [7]. Our contribution can be described as follows.
  • We formulated the closed-form gradient for the RoPE attention (see Lemma 5.3) along with its exact time complexity (see Theorem 5.7).
  • We derive the almost linear time backward approximation (see Theorem 6.7) for RoPE attention based on the closed-form gradient.
  • We show that with lower bounds derived from the SETH, the bounded entry condition is necessary for subquadratic performance (see Theorem 7.1).
@@
In Section 2, we present some relevant papers. In Section 3, we show essential components for the RoPE attention. In Section 4, we present the details of getting the gradients of RoPE-based attention mechanisms. Section 5 discusses the exact time complexity for computing RoPE attention gradients. Section 6 shows the fast computation of gradient using the polynomial method. Section 7 details the lower bounds of hardness. Finally, Section 8 provides conclusions and avenues for future work.
@@ 2 Related Work
In this paper, we study a variant of attention called RoPE attention. At a high level, RoPE gives more expressive power to the model in exchange for making the computational problem more complicated. In particular, many prior algorithms, such as the algorithm of [6], no longer apply to RoPE for fundamental reasons we will discuss. RoPE was proposed by [58] and has been used extensively in large-scale industrial models. Examples which are known to use RoPE include the open-source models released by Meta such as Llama [70] (see page 3), Llama 2 [71] (see page 5), Llama 3 [54] (see page 7), and the close-source LLM Claude 3.5 [4] released by Anthropic. Apple also incorporates RoPE into their LLM architecture (see [55], and page 3 of  [29]). The idea behind RoPE is to rotate the query and key vectors in the self-attention mechanism. The rotation is position-dependent and designed such that the inner product between two position-encoded vectors reflects their relative position in the input. In the RoPE attention mechanism, pairs of tokens with a longer relative distance will have a smaller correlation.
@@
The attention mechanism has often been criticized for its quadratic computational complexity concerning context length, a challenge that becomes more pronounced as the sequence length grows in today’s LLMs [2, 57, 54, 3, 5, 4]. However, this issue can be addressed using polynomial kernel approximation methods [1], which facilitate constructing the approximated attention matrix using low-rank approximations. Such techniques enable substantial improvements in computation speed, allowing a single attention layer to perform both training and inference nearly as fast as linear time [6, 8]. Our research further extends this efficiency to support multi-layer transformer architectures for both training and inference. In addition, these techniques can generalize to advanced attention mechanisms, such as tensor attention, while preserving the almost linear time complexity in both training and evaluation phases [9, 51]. Beyond this, alternative theoretical methods also exist. For example, the conv-basis approach introduced in [45] offers another avenue for speeding up attention computation.
@@
The attention has become a cornerstone in AI, particularly in large language models (LLMs), which excel in NLP tasks such as machine translation, text generation, and sentiment analysis due to their ability to capture complex contextual relationships. However, understanding the attention mechanism from a theoretical perspective remains an ongoing challenge. Several works have explored the theoretical foundations and computational complexities of attention [69, 80, 17, 6, 66, 18, 20, 43, 56, 68, 12, 22, 16, 65, 10], focusing on areas such as efficient attention [31, 59, 67, 40, 45, 51, 49, 47, 46, 21, 44, 35, 34, 30, 73, 32, 74, 36, 8, 24, 59, 41, 50, 60, 61], optimization [23, 52, 62, 27, 53, 26, 25], and the analysis of emergent abilities [15, 77, 11, 37, 79, 42, 78, 19, 48, 33, 76, 63, 75, 78, 38]. Notably, [80] introduced an algorithm with provable guarantees for attention approximation, [39] proved a lower bound for attention computation based on the Strong Exponential Time Hypothesis, and [6] provided both an algorithm and hardness results for static attention computation using low-rank approximation [28, 64].
@@ 3.1 Notation
For $n\in\mathbb{Z}^{+}\cup\{0\}$, for set $\{1,2,\cdots,n\}$, we denote the set by using the notation $[n]$. Here, we define the concept of nearly linear time when the time is $O(n\log n)$. We introduce the concept of almost linear time when time is $O(n^{1+o(1)})$. Given $a$ as any vector, we say the diagonal matrix of $c$ is $\operatorname{diag}(c)$ where $c_{i}$ means the $i,i$-th entry in the matrix $\operatorname{diag}(c)$. For any matrix, we denote the support of the matrix using the notation $\operatorname{supp}$, that is, the set of entries where the matrix is nonzero. $B^{\top}$ is defined as $(B^{\top})_{i,j}:=B_{j,i}$. Suppose there are two vectors $c,d$ of the same length. We denote the entry-wise multiplication using the notation $c\circ d$; that is, the $i$-th entry in that vector is $c_{i}d_{i}$. To denote the Frobenius norm, for any matrix $B$, we denote it as $\|B\|_{F}:=\sqrt{\sum_{i,j}B_{i,j}^{2}}$; to denote the maximum norm of matrix $B$, we use $\|B\|_{\infty}:=\max_{i,j}|B_{i,j}|$. Suppose there are two matrices $C,D$ of the same dimensions. We represent the Hadamard product or the entry-wise multiplication by using the notation $C\circ D$, that is, $(i,j)$-th entry of the matrix is $C_{i,j}\cdot D_{i,j}$. Let $C\in\mathbb{R}^{n_{0}\times m_{0}}$ and $D\in\mathbb{R}^{n_{1}\times m_{1}}$. We define $C\otimes D$ is an $n_{0}n_{1}\times m_{0}m_{1}$ matrix, where $(C\otimes D)_{(j_{0}-1)n_{1}+j_{1},(i_{0}-1)m_{2}+i_{1}}$ is equal to $C_{j_{0},i_{0}}D_{j_{1},i_{1}}$ for any $j_{0}\in[n_{0}],i_{0}\in[m_{0}],j_{1}\in[n_{1}],i_{1}\in[m_{1}]$.
@@ 3.2 Problem Definition
$A(\mathsf{X})_{i,j}=\exp(\underbrace{\mathsf{A}_{i+(j-1)n,*}}_{1\times d^{2}}\underbrace{\mathsf{X}}_{d^{2}\times d^{2}}\underbrace{w_{i-j}/d}_{d^{2}\times 1}).$
@@
$A(\mathsf{X})_{i,j}=$ $~{}\exp(\underbrace{(\mathsf{A}_{i+(j-1)n,*}\otimes w_{i-j}^{\top})}_{1\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1})=~{}\exp(\underbrace{(\mathsf{A}_{i+(j-1)n,*}\otimes\mathsf{W}_{i+(j-1)n,*})}_{1\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1}).$
@@
Then, the optimization problem in the context of RoPE attention computation is described as follows:
@@
Let $B>0$ and $\epsilon>0$ denote two parameters. Given a set of matrices $W_{-(n-1)},\cdots,W_{-1}$, $W_{0},W_{1},\cdots,W_{n-1}\in\mathbb{R}^{d\times d}$ where $\mathrm{supp}(W_{i})\subset S$ for all $i\in\{-(n-1),\cdots,-1,$ $0,1,\cdots,n-1\}$. Here $S\subseteq[d]\times[d]$ where $|S|=O(d)$. For $i,j\in[n]$, let $\mathsf{W}\in\mathbb{R}^{n^{2}\times d^{2}}$ such that $\mathsf{W}_{i+(j-1)n,*}=\operatorname{vec}({W_{i-j}})$. Here, we suppose four $n\times d$ matrices $A_{1},A_{2},A_{3},E$, and we have three $d\times d$ matrices $X_{1},X_{2},Y$. Let $\mathsf{X}:=X_{1}\otimes X_{2}\in\mathbb{R}^{d^{2}\times d^{2}}$. We define the matrix $A(\mathsf{X})\in\mathbb{R}^{n\times n}$ as the matrix representation of $\exp(\underbrace{(\mathsf{A}\oslash\mathsf{W})}_{n^{2}\times d^{4}}\underbrace{\operatorname{vec}({\mathsf{X}})/d}_{d^{4}\times 1})$ and the $n\times n$ matrix $D(\mathsf{X}):=\operatorname{diag}(\underbrace{A(\mathsf{X})}_{n\times n}\underbrace{{\bf 1}_{n}}_{n\times 1}).$ The RoPE attention optimization problem is formulated as follows:
@@ 3.4 Time Complexity of Multiplications
Matrix multiplication is a fundamental operation in many algorithms, and understanding its time complexity is essential for analyzing computational efficiency. Here, we introduce the time complexity of matrix multiplications.
@@
To further analyze the structure of matrix multiplication time complexity, we rely on a well-known fact from prior research [13, 14]. This fact provides equivalences between different permutations of matrix dimensions.
@@ 3.5 $\mathsf{SETH}$ Hypothesis
Now, we introduce a fundamental theoretical assumption underpinning many of the results presented in this paper: the Strong Exponential Time Hypothesis ($\mathsf{SETH}$). This hypothesis serves as a cornerstone for establishing the hardness of various computational problems.
@@ 3.6 Basic Facts
In this section, we present several basic facts that are used throughout the paper to develop the proof of our main results. These fundamental properties enable efficient computations of vectors and matrices products in our paper.
Here, we first introduce the facts about row-wise Kronecker products.
Fact 3.8 (Folklore).
@@
Next, we introduce some important properties of inner products that help us to reshape the equations in the proofs.
Fact 3.9 (Inner Products).
@@
To simplify the computation of certain matrix operations, we can use a technique known as the tensor trick, which reformulates matrix products into operations involving vectorized representations and Kronecker products.
Fact 3.10 (Tensor trick).
@@
Given the above tensor trick fact, we can derive additional properties that extend its applicability to exponential operations on matrices. These properties can help us compute the exponential of matrix products efficiently. The properties are presented below.
Fact 3.11.
@@
From Lemma and Def. 3.10, we are able to prove this fact. ∎
@@ 3.7 Useful Definitions
In this section, we decompose RoPE attention into its individual components, each representing a specific function or operation within the attention mechanism. These definitions provide a structured framework for understanding and analyzing the properties of RoPE attention in subsequent sections.
We denote the $d^{4}$-dimensional vector $x\in\mathbb{R}^{d^{4}}$ as the vectorization of a $d^{2}\times d^{2}$ matrix $\mathsf{X}$. We divide the RoPE attention to the following components to simplify our calculations and notation.
First, we define $u(x)$ for the softmax operation.
Definition 3.12 (Softmax $u(x)$).
@@
Next, we define $\alpha(x)$ for the diagonal matrix.
Definition 3.13 (Diagonal matrix $\alpha(x)$).
@@
Then, we define $s(x)$ for the normalized softmax $(D^{-1}\cdot\mathrm{softmax})$.
Definition 3.14 (Normalized softmax $s(x)$).
From Def. 3.12, it defines $u(\cdot)_{j_{0}}$ , and we have $\alpha(\cdot)_{j_{0}}$based on Def. 3.13. The function $s(x)_{j_{0}}$ maps a $d^{4}$-dimensional vector to an $n$-dimensional vector given every $j_{0}\in[n]$ such that
@@
Lastly, we define $v(y)$ for the value matrix in the attention component.
Definition 3.15 (Value matrix $v(y)$).
@@
Given the definitions of the RoPE attention components, we can now define the loss functions, which quantify the difference between the computed and target values in the context of RoPE attention.
We first introduce the error $\ell(x)_{j_{0},i_{0}}$ between the exact RoPE attention computation $\langle s(x)_{j_{0}},v(y)_{i_{0}}\rangle$ and approximated RoPE computation $E_{j_{0},i_{0}}$.
@@
From Def. 3.14, with every $j_{0}$ in the set $[n]$, it gives $s(x)_{j_{0}}$ as an $n$-dimensional normalized vector, and we define $v(y)_{i_{0}}$ based on Def. 3.15 given that each $i_{0}\in[d]$. Defining a function $\ell(x)_{j_{0},i_{0}}$ maps a $d^{4}$-dimensional vector to a scalar with each $j_{0}\in[n]$ and each $i_{0}\in[d]$ such that
@@ 3.8 Reformulation of the Loss Function
In this section, we are able to reformulate and simplify the loss function based on the definitions provided in Section 3.7. This reformulation provides a structured representation of the loss in terms of its components, using the tensor trick to simplify computations and facilitate analysis.
The following lemma formalizes this reformulation, consolidating the expressions for the loss function and connecting its components:
@@
  • Based on Def. 3.17, For $j_{0}$ in the set $[n]$ and $i_{0}$ in the set[d]$[d]$, we get $\mathsf{Loss}(\mathsf{X})_{j_{0},i_{0}}$
@@
where the 1st equality is based on Def. 3.2, the definition of Frobenius norm derives the 2nd equality, the 3rd equality is due to Def. 3.14 and Def. 3.15, and the 4th step is based on Def. 3.17. ∎